# Emergence of spin-phonon coupling in Gd-doped Y2CoMnO6 double perovskite oxide: a combined experimental and ab-initio study

### v2

@@ Abstract
One of the fundamental interactions that is found in many functional materials is the spin-phonon coupling (SPC), which is at the heart of many novel functionalities. The simultaneous presence of multi-magnetic phases makes SPC even more intriguing. We have used Raman spectroscopy as well as first-principles methods to investigate the possibility of the appearance of spin phonon coupling (SPC) in Gd-doped Y2CoMnO6 (YGCMO) double perovskite oxide and the influence of anti-site disorder on the same. YGCMO exhibits anti-site disorder leading to both ferromagnetic (between Co and Mn) and anti-ferromagnetic interactions (Co-Co, Mn-Mn, Gd-Co/Mn), making the SPC more intriguing. An analysis of the temperature-dependent phonon frequency for the stretching modes of YGCMO indicates that SPC is possibly emerging from simultaneous presence of ferromagnetic and antiferromagnetic interactions. The spin-phonon coupling strength comes out to be 0.29 cm-1. Our experimental findings are corroborated by first-principles density functional theory (DFT) based calculations which indicate that SPC in YGCMO gets enhanced in the presence of anti-site disorder. This implies a strong influence of B-site (Co/Mn) ordering on SPC in the bulk double perovskite systems. Thus, the phonon dynamics of YGCMO are strongly correlated with magnetic ordering. This material has potential applications in spintronics devices.
@@ 1 Introduction
In general, the ordered double perovskite (DP) R2BMnO6 materials are found to be ferromagnetic (FM) in nature by virtue of the B2+-O-Mn4+ superexchange interaction [6, 7]. Previous experimental and theoretical research on this class of materials reports the emergence of several novel functionalities, namely, magnetocapacitance, magnetoresistance, relaxor ferroelectricity, and multiferroicity [9, 10, 11, 12]. Besides, the nature of order/disorder at the B and Mn sites in these DP materials has been found to greatly influence the electronic, magnetic, vibrational, and dielectric properties [13, 14, 15, 5, 16]. Anti-site disorder (ASD) may also result in the suppression of ferromagnetism in some materials. This is because ASD induces B2+-O-B2+ and Mn4+-O-Mn4+ super-exchange interactions, which results in short range anti-ferromagnetic (AFM) interactions [28].
@@
Recently the DPs containing smaller rare earth cations have been of great research interest as many theoretical, as well as experimental studies report the emergence of novel multiferroic properties [18] in these materials. Of particular interest is the Y-based double perovskite such as Y2CoMnO6 (YCMO) and Y2NiMnO6, which have been found to exhibit polar nature in their magnetic ground state [19]. Moreover, first-principles calculations have shown that YCMO double perovskite oxides develop E* type magnetic ordering which is the source of ferroelectricity in these materials [20, 21]. It also exhibits some interesting properties like disorder-induced exchange bias and other multiferroic properties [28].
@@
In a strongly correlated system, the richness of the novel functionalities depends on the coupling of magnetic spins, the crystal lattice, and lattice vibrations. The interaction is complex but rich in the case of DP materials as it has two different magnetic species. So, a qualitative model based on spin-phonon coupling was proposed earlier for YCMO [22] in order to explain magnetically induced ferroelectricity in this material. SPC provides a fundamental background to understand various phenomena, such as magneto-thermal transport, magnetoelectric coupling, thermal Hall Effect, relaxation time in spintronics, etc. [23]. Spin-phonon coupling also facilitates the design of different spintronic devices with low-power and high-speed operations [int_S]. For instance, external magnetic fields are used to regulate magnetic refrigeration, isothermal entropy, and adiabatic temperature [int_S]. Furthermore, SPC also provides a chance to modulate the functionality of transition metal oxides (TMO) [25]. It is now well-understood that the coupling of phonons with other degrees of freedom may play a vital role in dictating several properties of these systems. In the past several studies have delved into the electronic, magnetic, optical, and transport properties of DP systems, phonon-related phenomena and their interactions with other degrees of freedom have not yet been thoroughly investigated. [5, 13]. Hence, an investigation of the nature of the “coupling” is crucial for understanding these systems better.
Interestingly, the spin- phonon coupling of DP system can be changed with lattice parameters such as cationic size, bond length, and bond angle [24]. Kumar et al. investigated that the SPC value gets weakened with the decrease in ionic radius of A-site cation in DP thin film [24]. This is related to the induced chemical pressure (large octahedral distortion) in the system, which modifies the FM superexchange interaction due to decrease in bond length of TM ions. On the other hand, Co-based DP compound LCMO exhibits SPC in thin films, whereas phonon modes remain unaffected due to change in magnetic ordering in bulk sample [26].
In this study, we have replaced 50% Y3+ ions with magnetic rare earth Gd3+ ions [28] to account for the effect of both local distortion and magnetism on SPC in YGCMO compound. For this purpose, temperature-dependent Raman spectroscopy along with first-principles DFT calculations are employed to investigate SPC in this DP compound [27]. It can also unveil the origin of SPC, whether it is direct or due to magnetostriction.
@@ 2 Experimental and Computational Details
The polycrystalline, single-phase double perovskite sample of Y1-xGdxCoMnO6 (=0.5) (YGCMO) was synthesized by the well-known chemical sol-gel method. A detailed description of sample preparation can be found in section S1 of the supplementary information (SI).
@@
The detailed structural and magnetic study on YGCMO has been done in our previous paper [28] which reveals that the compound crystallizes in monoclinic symmetry as a prime phase (99.42) [28]. It is partially disordered with Co2+, Co3+, Mn3+ and Mn4+ ions. An appreciable amount (38%) of antisite disorder is found in this sample. It exhibits three different magnetic transitions: i) FM transition triggers at TC= 95.5 K due to Co2+-O-Mn4+ superexchange interaction. ii) Antiferromagnetic (AFM) transition occurs at TN= 47 K due to exchange interaction of Co/Mn and Gd sublattices and ASD disorder present in the sample. iii) At T20 K, Gd3+-O-Gd3+ interactions become active and give rise to AFM interactions. More details are found in SI (refer to Fig.S1).
@@ 3 Results and discussion
The irreducible representation of the possible Raman modes of Gd-substituted YCMO at T=4 K is shown in Fig. 1(a). The absence of most intense Raman peak of Y2O3 near 377 cm-1 [40, 41] indicates the absence of a secondary phase (RE2O3) [42] in the system. Fig. 1(a) depicts eleven first-order Raman modes resolved at T=4 K, and their frequencies are determined by deconvoluting the spectrum with the Lorentzian line profile. However, no structural phase transition is observed throughout the temperature variation, confirming that the YGCMO remains in a monoclinic structure with space group , shown in Fig. 1 (b), (c), and (d). In this symmetry, 4e sites are generally occupied by the rare earth ions and oxygen atoms, whereas Co and Mn ions prefer 2c and 2b sites, respectively. The schematic of the structure is shown in Fig. 5.
The low symmetry structures of double perovskites result from a distorted cubic Fmm aristotype lattice [43]. From the group theory calculation, 24 Raman active modes are expected to be observed for this space group. Its distribution according to the irreducible representation of 2/m factor point group is given as follows: , which can be written as 12Ag+12Bg [27]. Lattice modes (= translational and = librational) are observed below 380 cm-1 whereas internal modes corresponding to oxygen octahedron ( and ) are detected above 380 cm-1 [27]. Generally, vibrational modes correspond to oxygen atoms connected with higher valence ions [44]. In this experiment, out of 24 Raman active modes, 11 Raman modes can be resolved at T= 4 K which is notably larger than the resolved mode for other La-based double perovskites [15, 27]. Smaller ionic radii of Y3+ and Gd3+ cation (r = 1.20 Å and r = 1.22 Å) in YGCMO lead to more distortion in the structure, which results in major octahedra tilts. As a consequence, a significant number of modes have been observed. In comparison with LCMO, the Raman bands split more for YGCMO due to strong monoclinic distortion [45].
@@
In the past, Iliev et al. [27] have identified the Raman modes for Mn-based DP system from the semiempirical calculation, and it is in good agreement with their experimental results. Therefore, based on some of the previous works [27, 46], Raman modes at 632 cm-1, 505 cm-1, 340 cm-1, 298 cm-1, 265 cm-1, and 196 cm-1 observed in our experiment can be assigned to Ag symmetry and also known as stretching mode (S). Here Co/Mn-O bond involves both in-plane and out-of-plane vibrations and those detected at 544 cm-1, 476 cm-1, and 406 cm-1 have Bg symmetry and are also known as anti-stretching mode (AS). Here O2- ions vibrate perpendicularly to the Co/Mn-O bond. The mode noticed at higher wave number 1270 cm-1 corresponds to the second-order overtone of breathing mode [47].
In order to study spin-phonon coupling in our sample, the effect of temperature on the phonon modes has been analyzed. The recorded temperature dependence of the Raman spectrum between 4 K and 300 K is shown in Fig. 1 (b), (c), and (d). These figures illustrate the impact of magnetic ordering on the lattice vibration. Generally, in manganite perovskite, the stretching mode shows spin-phonon coupling [48]. In our study, the temperature dependence of two of the most representative Raman modes are presented in Fig. 2 (a-b). Temperature dependence of phonon frequency due to anharmonic contribution (due to lattice contraction) proposed by Balkanski [49] can be written as follows,
@@
where (anharmonic constant) and (position of the Raman mode at the temperature T=0 K) are the adjustable parameters. In the absence of any phase transition, temperature-dependent Raman mode should follow the above anharmonic relation. It is seen that the most intense Raman mode, also known as stretching mode (observed at 632 cm-1), fits nicely with the anharmonic model above , shown in Fig. 2(a). This mode shows a deviation from the above model for TTC with two distinct discontinuities at TC and TN, respectively. The deviation is more pronounced with lowering the temperature. This anomalous behavior could arise due to the renormalization of phonons induced by the ferromagnetic ordering of Co-Mn ions and results in the coupling between spin and lattice (phonon) degrees of freedom [50]. A similar feature is observed in several double perovskite manganite systems [51, 52]. On the other hand, the AS Raman mode observed at 406 cm-1, fits well with the above anharmonic model throughout the investigated temperature range except near TC. At lower temperatures, the phonon frequencies become nearly constant for this mode. This temperature-independent behavior of phonon frequencies at lower temperatures could happen due to a change in magnetic spin structure or weak ferromagnetic ordering [53]. Our previous study shows that below TGd (Gd ordering temperature), Gd spins are at ab plane, which was in direction [28].
@@
The temperature dependence of linewidth of stretching and anti-stretching modes are plotted in Fig. 3 (a-b) to investigate the role of magnetostriction in this coupling. Raman linewidth is associated with phonon life-time that is not affected by lattice volume changes i.e. magnetostriction effect. Generally, a significant anharmonic effect has been observed at higher temperatures due to rigorous atomic vibration of the atoms about their mean position. However, this anharmonic effect gets reduced with a decrease in temperature due to reduced phonon-phonon interactions. As a result, phonon-lifetime increases or FWHM decreases. We observed that both the curves show anomaly upon entering into the magnetic region with the discontinuities at 95.5 K () and 47 K (), respectively, which suggests relative dependence of phonon lifetime with magnetic ordering. Therefore, strong interaction between phonon and magnetic degree of freedom leads to an increase in linewidth at lower temperatures i.e., a decrease in phonon lifetime. Usually, the linewidth is affected by spin-phonon coupling and electron-phonon coupling [59]. The total electronic density of states for YGCMO shows that it is insulating in nature(refer to Fig.S2 in SI). Therefore, the anomaly in linewidth with temperature is ascribed to spin-phonon coupling [59].
The careful observation of the above temperature-dependent frequency and linewidth plot reveals a new phonon renormalization arising at TN due to AFM ordering of Gd-Co/Mn spins, implying the presence of spin-phonon coupling. This result suggests that the FM phase i.e. ordered regions and similar disordered clusters exist together in this sample, where the predominant AFM interactions are produced from Co3+-O- Co3+, Mn3+-O-Mn3+, Co2+-O- Co2+, Mn4+-O-Mn4+, and - exchange pathways. Besides magnetic measurements, Raman spectroscopy is a powerful tool to detect magnetic inhomogeneity present in the sample.
@@
To get detailed insight into the anomalous behavior of Raman modes, the renormalization of phonon frequency with normalized magnetic moment has been plotted in Fig. 4(a). Usually, the deviation of phonon mode from the anharmonic model can be explained by the spin-spin correlation function in magnetic oxides [26].
@@
where is the normalized phonon frequency due to SPC, is the phonon frequency in the absence of SPC, and are nearest neighbor spins at and sites, respectively. denotes the strength of spin-phonon coupling. is obtained by taking the difference between calculated and observed phonon frequencies. According to mean-field theory, the spin-spin correlation function can be connected to magnetization by the equation , where and are the magnetization at T and saturation magnetization approximated to T= 0 K, respectively. Therefore, by considering the four nearest neighbors for each B site cation, can be written as follows [56]:
@@
The spin-phonon coupling constant value is estimated by using the above approach. Its value can be positive or negative depending upon phonon hardening or softening. In Fig. 4(a), it is seen that there is a good overlap of both and normalized magnetization up to 95 K, after that a deviation in the magnitude of occurs. It signifies that the origin of anomalous softening is ascribed to spin-phonon coupling. The occurrence of this unconventional behavior below 95 K might be caused by competition between FM and AFM interactions due to ASD present in the sample. ASD helps to originate different types of AFM (Co2+-O-Co2+, Mn4+-O-Mn4+, etc.) along with FM (Co2+-O-Mn4+) interactions, leading to magnetic frustration in the systems. In our sample, another source of AFM is the interaction between Co/Mn and Gd sublattices. Therefore, different values of magnetic coupling can contribute in different ways to the phonon renormalization induced by magnetic ordering. It makes the spin-phonon coupling more complex and causes deviation from mean field approximation. A similar kind of deviation is observed in La2CoMnO6 [55], Gd2CoMnO6 [56], and Lu2CoMnO6 [57], where FM and AFM interactions coexist together.
Spin-phonon coupling strength () is determined by taking a quarter of the slope of the linear relationship of vs. [24] shown in Fig. 4 (b). The value is found to be 0.29 cm-1. It has been observed that SPC value is always lower for bulk sample than thin film [26] because epitaxy stress plays pivotal role towards SPC. Silva et al. reveal that ordered bulk LCMO does not show SPC whereas disordered GCMO exhibits SPC to a greater extent [60].
Again, Meyer et al. [26] show that spin-phonon coupling coefficient value is lower for disordered LCMO in thin film than the ordered one. Thus B-site cationic ordering has not only influenced the magnetic properties of YGCMO perovskite significantly but also affects its phonon properties and spin-phonon coupling strength to a great extent.
@@
SPC in the YGCMO compound is further investigated by first-principles calculations. We have considered a supercell of size 121 containing 40 atoms i.e, there are 4 Y, 4 Gd, 4 Co, 4 Mn, and 24 O atoms in each supercell. The structure was fully optimized, starting from the room temperature experimental data. Fig. 5(a) shows the lowest energy ordered structure of YGCMO where Co and Mn octahedra are arranged alternately. To determine the influence of anti-site disorder on the nature of SPC, starting from the ordered 121 crystal structure, we have manually introduced some anti-site disorder. This disordered structure thereafter underwent a complete optimization (see Fig. 5 (b)). The total electronic density of states for the disordered structure (which contains Co-Co, Co-Mn, and Mn-Mn interactions) shows that the sample is insulating with a band gap of 1 eV (refer to Figure S2 of the SI).
@@
To determine the magnetic ground state, we have performed total energy calculations on the ferromagnetic, C-, G-, and A-type antiferromagnetic and ferrimagnetic configurations of ordered-YGCMO. These magnetic interactions were considered among the Co and Mn ions, whereas the Gd moments were made to orient along the z-direction in all the calculations. Our calculations indicate that the ferromagnetic state (see Fig. 6(a)) has the lowest energy followed by the C-type AFM magnetic configuration (see Fig. 6 Y(b)). The relative energies of the different magnetic orderings with respect to the ferromagnetic state have been tabulated in Table 1.

### v3

# Emergence of spin-phonon coupling in a Gd-doped Y2CoMnO6 double perovskite oxide: a combined experimental and ab-initio study

@@ Abstract
We present Raman spectroscopy results backed by first-principles calculations and investigate the nature of possible spin-phonon coupling (SPC) in a Gd-doped Y2CoMnO6 (YGCMO) double perovskite oxide. The influence of Gd substitution, A-site ordering, and anti-site disorder is also studied. YGCMO exhibits anti-site disorder leading to both ferromagnetic (between Co and Mn) and antiferromagnetic interactions (Co-Co, Mn-Mn, Gd-Co/Mn), making the SPC quite intriguing. An analysis of the temperature-dependent phonon frequencies for the stretching modes of YGCMO indicates that SPC here possibly emerges from the simultaneous presence of competing ferromagnetic and antiferromagnetic interactions. The SPC strength comes out to be 0.29 cm-1. Our density functional theory (DFT) calculations show that Phonon modes shifted towards lower frequency with Gd doping. Similarly, A-site ordring and anti-site disorder significantly alter the Raman spectra. Experimental findings are also corroborated by first-principles DFT calculations, which indicate that anti-site disorder and Gd doping enhances SPC in YGCMO. This implies a strong influence of A-site cationic radii, and B-site (Co/Mn) ordering on SPC in the bulk double perovskite systems. The phonon dynamics of YGCMO are, therefore, correlated with magnetic ordering, indicating potential applications in spintronics devices.
@@ 1 Introduction
In general, the ordered double perovskite (DP) R2BMnO6 materials are found to be ferromagnetic (FM) in nature by virtue of the B2+-O-Mn4+ superexchange interaction [6, 7]. Previous experimental and theoretical research on this class of materials reports the emergence of several novel functionalities, namely, magnetocapacitance, magnetoresistance, relaxor ferroelectricity, and multiferroicity [9, 10, 11, 12]. Besides, the nature of order/disorder at the B and Mn sites in these DP materials has been found to greatly influence the electronic, magnetic, vibrational, and dielectric properties [13, 14, 15, 5, 16]. Anti-site disorder (ASD) may also result in the suppression of ferromagnetism in some materials. This is because ASD induces B2+-O-B2+ and Mn4+-O-Mn4+ super-exchange interactions, which result in short-range anti-ferromagnetic (AFM) interactions [31].
@@
Recently, the DPs containing smaller rare earth cations have been of great research interest as many theoretical, as well as experimental studies report the emergence of novel multiferroic properties [18] in these materials. Of particular interest is the Y-based double perovskite such as Y2CoMnO6 (YCMO) and Y2NiMnO6, which have been found to exhibit polar nature in their magnetic ground state [19]. Moreover, first-principles calculations have shown that YCMO double perovskite oxides develop E* type magnetic ordering which is the source of ferroelectricity in these materials [20, 21]. It also exhibits some interesting properties like disorder-induced exchange bias and other multiferroic properties [31].
@@
In a strongly correlated system, the richness of the novel functionalities depends on the coupling of magnetic spins, the crystal lattice, and lattice vibrations. The interaction is complex but rich in the case of DP materials as it has two different magnetic species. So, a qualitative model based on spin-phonon coupling was proposed earlier for YCMO [22] in order to explain magnetically induced ferroelectricity in this material. Moreover, SPC provides a fundamental background to understand various phenomena, such as magneto-thermal transport, magnetoelectric coupling, thermal Hall Effect, relaxation time in spintronics, etc. [23, 24]. Spin-phonon coupling also facilitates the design of different spintronic devices with low-power and high-speed operations [28]. For instance, external magnetic fields are used to regulate magnetic refrigeration, isothermal entropy, and adiabatic temperature [28]. Furthermore, SPC also provides a chance to modulate the functionality of transition metal oxides (TMO) [29]. It is now well-understood that the coupling of phonons with other degrees of freedom may play a vital role in dictating several properties of these systems. We see that several studies have delved into the electronic, magnetic, optical, and transport properties of DP systems, but very few phonon-related phenomena and their interactions with other degrees of freedom have been investigated in their bulk form. It has been reported that a large spin-phonon coupling is observed in thin films of A2CoMnO6 (A = rare earth) double perovskite (DP) compounds, evidenced by significant softening of phonon modes below the Curie temperature (TC), which is absent in their bulk counterparts [25]. Similarly, Raman spectroscopy on La2CoMnO6 revealed subtle shifts in phonon frequency but substantial changes in linewidth across TC, attributed to lattice effects [26]. Kumar et al. further investigated that the softening of the phonon mode decreases with decreasing ionic radius of A-site cation in bulk DP compounds [25]. Conversely, Macedo Filho et al. found that in RE2NiMnO6 (RE = Pr, Nd, Gd, Tb, Y) compounds, the strength of spin-phonon coupling is largely independent of the A-site cation’s ionic radius and is primarily governed by B-site cationic ordering [27]. Hence, investigating the nature of the “coupling” is crucial for a better understanding of these systems.
In our previous work [31], we reported the experimental and theoretical investigation of the structural, electronic, and magnetic properties of 50% Gd-substituted YCMO. 50% doping at the rare-earth site helps to maintain a balanced charge state and structural stability of the doped compound. The substitution of larger Gd ions at A-site induces structural distortion, which might be coupled with the spin state of YGCMO. So, in this study, we explore the spin-phonon coupling (SPC) in YGCMO and the roles of Gd substitution and anti-site disorder on it using temperature-dependent Raman spectroscopy (serves as a powerful tool to investigate ordering, magnetism, and spin-phonon interactions) [32] and first-principles density functional theory (DFT) calculations. This approach enables us to elucidate the origin of SPC, distinguishing between direct coupling and magnetostriction effects.
@@ 2 Experimental and Computational Details
The polycrystalline, single-phase double perovskite sample of Y2-xGdxCoMnO6 (=1.0) (YGCMO) was synthesized by the well-known chemical sol-gel method. A detailed description of sample preparation can be found in Ref. [31]
@@ 3 Results and discussion
The detailed study on structure, oxidation states, and the magnetic states of the YGCMO compound has been reported earlier [31]. The crystal structure belongs to the space group (refer to Fig.S1 within the SI). In this symmetry, 4e sites are generally occupied by the rare earth ions (Y/Gd) and oxygen atoms (O), whereas Co and Mn ions prefer 2c and 2b sites, respectively. The schematic of the structure is shown in Fig. 1 (a). The XPS study has shown that transition metal ions Co and Mn are present in their mixed valence states (refer to Fig.S2 within the SI), i.e., Co2+, Co3+, Mn3+, and Mn4+, respectively. The compound is found to be partially disordered with Co-Mn ions, and the calculated value of disorder becomes 38%. Finally, the magnetic measurement reveals that the compound exhibits three different magnetic transitions: i) Transition metal ions (Co/Mn) mediated FM transition triggers at TC= 95.5 K. ii) Antiferromagnetic (AFM) transition occurs at TN= 47 K due to exchange interaction of Co/Mn and Gd sublattices and ASD disorder present in the sample. iii) At T20 K, Gd3+-O-Gd3+ interactions become active and give rise to AFM interactions. More details are found in the supplementary information (refer to Fig.S3).
The Raman spectrum of 50% Gd-substituted YCMO recorded at T=4 K within the spectral range between 100 and 800 cm-1 is shown in Fig. 1(b). The absence of the most intense Raman peak of Y2O3 near 377 cm-1 [44, 45] discards the presence of a secondary phase (RE2O3) [46] in the system. Generally, the low symmetry structures with space group of double perovskites result from a distorted cubic Fmm aristotype lattice [47]. The group theory analysis predicted 60 first-order gamma-point phonon modes, of which 24 correspond to Raman active modes and 36 are infrared active modes. Its distribution according to the irreducible representation of 2/m factor point group is given as follows: , which can be written as 12Ag+12Bg [32], where lattice modes (= translational and = librational) and internal modes corresponding to oxygen octahedron ( and ) are observed below and above 380 cm-1 [32]. We observed eleven first-order Raman modes below 800 cm-1 and their frequencies are determined by deconvoluting the spectrum with the Lorentzian line profile. A significant number of deconvoluted modes have been observed due to more structural distortion, primarily caused by the substitution of Gd3+ (r = 1.22 Å) cation at Y3+ (r = 1.20 Å) site. The mode noticed at higher wave number 1270 cm-1 corresponds to the second-order overtone of breathing mode shown in Fig. 2 (c) [52]. Later, we subsequently explore the influence of Gd doping and Gd ordering on the Raman spectrum, alongside the impact of antisite disorder (ASD) on the Raman spectrum, through theoretical studies.
@@
In the past, Iliev et al. [32] have identified the Raman modes for Mn-based DP system from the semiempirical calculation, and it is in good agreement with their experimental results. Therefore, based on some of the previous works [32, 51], Raman modes at 632 cm-1, 505 cm-1, 340 cm-1, 298 cm-1, 265 cm-1, and 196 cm-1 observed in our experiment can be assigned to Ag symmetry and also known as stretching mode (S). Here Co/Mn-O bond involves both in-plane and out-of-plane vibrations and those detected at 544 cm-1, 476 cm-1, and 406 cm-1 have Bg symmetry and are also known as anti-stretching mode (AS). Here O2- ions vibrate perpendicularly to the Co/Mn-O bond. Generally, vibrational modes correspond to oxygen atoms connected with higher valence ions [48].
Now, we focus on some first-order Raman modes and the evolution of their Raman shifts with temperature. The recorded temperature dependence of the Raman spectrum between 4 K and 300 K is shown in Fig. 2 (a), (b), and (c). According to, Granado et al.[49], the change in phonon frequency with temperature in magnetic materials can be expressed as follows: , where is the phonon frequency at 0 K. corresponds to the change in phonon wave-number due to the thermal expansion of lattice. The change in phonon frequency due to intrinsic anharmonic vibrational potential at constant volume is given by . and are the spin-phonon and electron-phonon coupling terms, respectively.
can be determined by the change in unit-cell volume with temperature. However, no spectral changes are observed throughout the temperature variation shown in Fig. 2 (a), (b), and (c), confirming that the YGCMO remains in a monoclinic structure with space group . Therefore, the observed phonon frequency shift is probably not caused by the lattice expansion with temperature.
The anharmonicity in lattice vibration can modify the Raman shift () and FWHM of the phonon modes at higher temperature. In general, the temperature dependence of phonon frequency due to anharmonic contribution is proposed by Balkanski [54] can be written as follows,
@@
where (anharmonic constant) and (position of the Raman mode at the temperature T=0 K) are the adjustable parameters. In our study, the temperature dependence of phonon frequency for two of the most representative Raman modes is analyzed using Eq. 1 and presented in Fig. 3 (a-b). In the absence of any phase transition, temperature-dependent Raman mode should follow the above anharmonic relation. It is seen that the most intense Raman mode, also known as stretching mode (observed at 632 cm-1), fits nicely with the anharmonic model above i.e, the compound is in the paramagnetic region, shown in Fig. 3(a). This mode shows a deviation from the above model for TTC with two distinct discontinuities at TC and TN, respectively. The deviation is more pronounced at lower temperature, which implies that the phonon frequency is affected by many other factors, apart from the anharmonicity. This anomalous behavior could arise due to the renormalization of phonons induced by the ferromagnetic ordering of Co-Mn ions and results in the coupling between spin and lattice (phonon) degrees of freedom [55]. A similar feature is observed in several double perovskite manganite systems [56, 57]. On the other hand, the AS Raman mode observed at 406 cm-1, fits well with the above anharmonic model throughout the investigated temperature range except near TC. At lower temperatures, the phonon frequencies become nearly constant for this mode. This temperature-independent behavior of phonon frequencies at lower temperatures could happen due to a change in magnetic spin structure or weak ferromagnetic ordering [58]. Our previous study shows that below TGd (Gd ordering temperature), Gd spins are at ab plane, which was in any other direction [31].
@@
Now, we investigate the role of magnetostriction on the observed phonon anomaly. In this regard, the study of FWHM or Raman linewidth is more effective than the position with temperature. Raman linewidth is associated with phonon life-time that is not affected by lattice volume changes, i.e., magnetostriction effect. Moreover, FWHM is affected by magnetic transition and reveals an anomaly in the FWHM near the transition. Therefore, FWHM is more effective for determining whether the anomaly arises from direct spin-phonon coupling or is induced by magnetostriction. Fig. 4 (a-b) shows the temperature dependence of the linewidth of the showing anomaly by stretching and anti-stretching modes of YGCMO. Generally, a significant anharmonic effect has been observed at higher temperatures due to the rigorous atomic vibration of the atoms about their mean position. However, this anharmonic effect is reduced with a decrease in temperature due to a decrease in phonon-phonon interactions. As a result, the phonon lifetime increases or the FWHM decreases. We observed that both curves show anomaly upon entering into the magnetic region with the discontinuities at 95.5 K () and 47 K (), respectively, suggesting the relative dependence of phonon lifetime with magnetic ordering. Therefore, strong interaction between phonon and magnetic degree of freedom leads to an increase in linewidth at lower temperatures i.e., a decrease in phonon lifetime. Usually, the linewidth is affected by spin-phonon coupling and electron-phonon coupling [64]. The total electronic density of states for YGCMO shows that it is half-metallic (refer to Fig.S5 (b) in SI). Therefore, the anomaly in linewidth with temperature is ascribed to spin-phonon coupling [64]. The careful observation of the above temperature-dependent frequency and linewidth plot reveals a new phonon renormalization arising at TN due to AFM ordering of Gd-Co/Mn spins, implying the presence of spin-phonon coupling.
@@
To get detailed insight into the anomalous behavior of Raman modes, the renormalization of phonon frequency with normalized magnetic moment has been plotted in Fig. 5(a). Usually for magnetic oxide, the deviation of phonon mode from the anharmonic model can be explained by the spin-spin correlation function in magnetic oxides [30].
@@
where is the renormalized phonon frequency due to SPC, is the phonon frequency in the absence of SPC, and are nearest neighbor spins at and sites and is the statistical average of the spin pair correlation function, respectively. denotes the strength of spin-phonon coupling. is obtained by taking the difference between calculated and observed phonon frequencies. According to mean-field theory, the spin-spin correlation function can be connected to magnetization by the equation , where and are the magnetization at T and maximum magnetization, respectively. Therefore, by considering the four nearest neighbors for each B site cation, can be written as follows [61]:
@@
The spin-phonon coupling constant value is estimated by using the above approach. Its value can be positive or negative depending upon phonon hardening or softening. Fig. 5(a) depicts that there is a good overlap of both and normalized magnetization near the transition temperature, beyond which a deviation in the magnitude of occurs. It signifies that the origin of anomalous softening is ascribed to spin-phonon coupling. The occurrence of this unconventional behavior below 95 K might be caused by competition between FM and AFM interactions. In this sample, the ferromagnetic (FM) phases coexist with antiferromagnetic (AFM) phases, where the predominant AFM interactions, such as Co3+-O- Co3+, Mn3+-O-Mn3+, Co2+-O- Co2+, and Mn4+-O-Mn4+ exchange pathways are produced due to ASD. Additionally, Gd-4f and Co/Mn-3d interaction is present in the sample. Thus, the magnetic frustration emerges from the competition between FM and AFM interactions.
Therefore, different values of magnetic coupling can contribute in different ways to the phonon renormalization induced by magnetic ordering. It makes the spin-phonon coupling more complex and causes deviations from the mean-field approximation. A similar kind of deviation is observed in La2CoMnO6 [60], Gd2CoMnO6 [61], and Lu2CoMnO6 [62], where FM and AFM interactions coexist together.
Spin-phonon coupling strength () is determined by taking a quarter of the slope of the linear relationship of vs. [25] shown in Fig. 5 (b). The value is found to be 0.29 cm-1. The spin-phonon coupling (SPC) strength is notably weaker in bulk systems compared to thin films, with a reported SPC value of = 0.51 for bulk Pr2CoMnO6 [25]. Silva et al. reported that ordered bulk LCMO does not show SPC, whereas disordered Gd(Co1/2Mn1/2)O3 exhibits a significant SPC [65]. Therefore, our sample, containing 38% ASD as stated above, shows SPC, and DFT calculations further reveal that SPC is enhanced in the presence of ASD. Thus, B-site cationic ordering in YGCMO perovskite profoundly influences not only its magnetic properties but also its phonon dynamics and spin-phonon coupling strength .
@@
The impact of Gd doping, Gd ordering, and the impact of ASD on Raman spectra and SPC in the YGCMO compound is further investigated by first-principles calculations. We have considered a supercell of size 121 containing 40 atoms of YCMO and YGCMO i.e, there are (8 Y, 4 Co, 4 Mn, 24 O) and (4 Y, 4 Gd, 4 Co, 4 Mn, and 24 O) atoms in each supercell. The structure was fully optimized, starting from the room temperature experimental data. Fig. 6(a) shows the lowest energy ordered structure of YGCMO where Co and Mn octahedra are arranged alternately. To determine the influence of anti-site disorder on the nature of SPC, starting from the ordered 121 crystal structure, we have manually introduced some anti-site disorder. This disordered structure thereafter underwent a complete optimization (see Fig. 6 (b)). The total electronic density of states for the disordered structure (which contains Co-Co, Co-Mn, and Mn-Mn interactions) shows that the sample is half-metal with a band gap of 1.1 eV in down spin channel(refer to Figure S5 (b) of the SI).
@@
To determine the magnetic ground state, we have performed total energy calculations on the ferromagnetic, C-, G-, and A-type antiferromagnetic and ferrimagnetic configurations of ordered and disordered-YGCMO. These magnetic interactions were considered among the Co and Mn ions, whereas the Gd moments were made to orient along the z-direction in all the calculations. Our calculations indicate that the ferromagnetic state (see Fig. 7(a)) has the lowest energy, followed by the C-type AFM magnetic configuration (see Fig. 7 (b)). The relative energies of the different magnetic orderings with respect to the ferromagnetic state have been tabulated in Table 1.